An important amino acid has the percentage composition C 32.00%, H 6.71%, and N 18.66%.
Calculate the empirical formula of this substance.
1
Expert's answer
2020-10-21T12:51:58-0400

If we consider the total mass of hydrocarbon is 100 gram, atomic mass of carbon is 12 gram/ mole, number of moles of carbon present in the hydrocarbon is 32/12=2.67mole , it is nearly equal to 3 mol,atomic mass of hydrogen is 1 gram/mole, number of moles of hydrogen present is 6.72/1=6.71mole, it is nearly equals to 7:00 mole, atomic mass of nitrogen is 14 gram per mole, number of moles of nitrogen present is 18.66/14=1.33, it is nearly equals to 1 mole,hence the empirical formula of the hydrocarbon is C3H7N